OpenStudy (anonymous):

a number dividwhened by 143 leaves 31 as remainder what will be the remainder when the same number is divided by 13

OpenStudy (anonymous):

A number when divided by 143 leaves 31 as remainder. ⇒ The given number is of the form 143x + 31 = 13(11x) + 13 × 2 + 5 = 13(11x + 2) + 5 ⇒ The remainder when the given number is divided by 13 is 5

OpenStudy (anonymous):

copy and pasted^^^^

OpenStudy (anonymous):

can u please explain it